OpenStudy (anonymous):

Kathryn currently has an account balance of $2,248.95. She opened the account 7 years ago with a deposit of $1,762.14. If the interest compounds quarterly, what is the interest rate on the account?

OpenStudy (anonymous):

interest=2248.95-1762.14=486.81 so interest= 486.81=1762.14*x/100 x=27.62

OpenStudy (anonymous):

wait are you trying to find x? cuz i dont think that is the interest rate. the interest rate would be a percentage, right?

OpenStudy (anonymous):

Yes it would be a percentage

OpenStudy (anonymous):

so 27.62%

OpenStudy (anonymous):

because the intestest rate would be like ending amount minus starting amount divided by 7 years divided by 4(because theres 4 quarters in each year) and then we get a number. and whatever percentage that is from the starting amount is the annual percentage i think?

OpenStudy (anonymous):

that doesnt equal the ending amount though.. because it's been 7 years and 4 quarters each year so 7*4 is 28 right? so according to your solved annual rate, we would multiply 28 * .2782 (because thats 27.82% as a decimal) and that result PLUS 1762.14 should give you 2248.95 but it doesn't..

OpenStudy (anonymous):

The choices are:0.9%,3.5%,9.0%,and 14.2%

OpenStudy (anonymous):

yes